A curve along the barangay is to be designed for motorela moving at 30 kilometers per hour. Find



the angle of banking if the radius of the curve is 20 m.

1
Expert's answer
2022-01-26T10:31:23-0500
"\theta=\arctan\dfrac{v^2}{rg}=\arctan\dfrac{30·10^3/3600}{20·9.8}=2.43°."